MAGNETIC FIELD TUNING OF THE NO_{2} PHOTODISSOCIATION THRESHOLD

\underline{REMY JOST}, ANTOINE DELON, JESPER NYGÅRD AND ADAM PASINSKI, Grenoble High Magnetic Field Laboratory, CNRS, BP 166, 38042 Grenoble Cedex 9, France.

The lowest photodissociation threshold of NO_{2}, located at 25128.57±0.05 cm^{−1}, is now well characterized: below the threshold, the highly excited vibronic levels of NO_{2}, observed by LIF in a jet, have a lifetime of a few μsec; above the threshold, the resonant levels have a lifetime shorter than 10^{2} psec and no fluorescence can be observed, because the molecule dissociates. The products, NO and O, are formed in their ground states: NO_{2}(X^{2}A_{1}) + hν → NO(^{2}Π_{1/2}) + O(^{3}P_{2}). We have studied by LIF the effect of a magnetic field on the dissociation threshold. The energy of the ground state of NO_{2} (N=0, K=0, J=1/2) is lowered by −μ_{B}B for the M_{S}=1/2 component. The energy of the NO is unaffected (g≈0 for the ^{2}Π_{1/2} state), while the energy of the oxygen atom is splitted into 5 components (J=2). The lowest component, M_{J}=−2, which correspond to the energetic threshold, is lowered by −3μ_{B}B, because the ^{3}P_{2} state has a Landé factor of 3/2. Consequently, we predict that the energy of the photodissociation threshold is lowered by −2μ_{B}B. We have experimentally and quantitatively observed this effect: the photodissociation threshold observed by LIF is lowered by 13.2 cm^{−1} in a magnetic field of 14 Tesla! The fact that we have to consider only the energy of the upper levels (and not their quantum numbers) is a consequence of the strong mixing (chaos) within the rovibronic levels of NO_{2} near the dissociation energy^{a}.

aR.Georges, A.Delon and R.Jost, J.Chem.Phys. 103, 1732 (1995). A.Delon, R.Georges and R.Jost, J.Chem.Phys. 103, 7740